# Robotic Learning for Informative Path Planning

### v1

@@ 3.3 Benchmarks
Simulators for AIPP evaluation make various interdependent design choices leading to highly specific simulation environments. First, AIPP methods are typically evaluated on a single fixed task, e.g., the RockSample task (Ott et al., 2023; Choudhury et al., 2020), monitoring missions Popović et al. (2020); Rückin et al. (2022a); Westheider et al. (2023); Cao et al. (2023a); Choi and Cielniak (2021); Hollinger and Sukhatme (2014), or reconstruction (Zeng et al., 2022; Maboudi et al., 2023; Mendoza et al., 2020; Gazani et al., 2023; Dhami et al., 2023; Ran et al., 2022; Zhan et al., 2023; Yan et al., 2023). Second, simulators differ in the robot configuration defined by the robot type, e.g., autonomous underwater vehicle (AUV(Duecker et al., 2021), unmanned aerial vehicle (UAV(Popović et al., 2020), unmanned ground vehicle (UGV(Wei and Zheng, 2020), or a robot arm (Zeng et al., 2022), as well as the sensors, e.g., depth sensors (Cao et al., 2023b; Liu et al., 2023; Niroui et al., 2019; Hepp et al., 2018), thermal camera (Rückin et al., 2022a), point-based sensors Cao et al. (2023a); Hitz et al. (2017); Yang et al. (2023a); Marchant et al. (2014), or abstractions of sensor models interpreting raw sensor data (Popović et al., 2020; Chaplot et al., 2021; Denniston et al., 2023a; Georgakis et al., 2022b; Denniston et al., 2023b). Third, works vary in their assumptions about the structure of environments, e.g., 2D terrains (Hitz et al., 2017; Popović et al., 2020; Cao et al., 2023a; Rückin et al., 2022a; Ott et al., 2023; Choudhury et al., 2020) or 3D volumes (Cao et al., 2023b; Zeng et al., 2022; Rayas Fernández et al., 2022), and robot workspaces, e.g., obstacle-free 2D (Cao et al., 2023a; Velasco et al., 2020), obstacle-free 3D (Popović et al., 2020; Westheider et al., 2023; Rückin et al., 2022a; Denniston et al., 2023b; Rayas Fernández et al., 2022), or obstacle-aware 3D (Cao et al., 2023b; Zeng et al., 2022; Hepp et al., 2018) workspaces. Last, simulators vary according to the datasets they use, e.g., synthetic environments or randomly generated distributions (Popović et al., 2020; Cao et al., 2023a; Hitz et al., 2017; Rückin et al., 2022a; Westheider et al., 2023; Schmid et al., 2022; Bai et al., 2017; Saroya et al., 2020), more realistic simulators (Zeng et al., 2022; Rückin et al., 2023; Zacchini et al., 2023; Zhang et al., 2023), or real-world remote sensing datasets Popović et al. (2020); Rückin et al. (2023); Westheider et al. (2023); Rayas Fernández et al. (2022); Hitz et al. (2017); Hollinger and Sukhatme (2014); Caley et al. (2019). Sec. 4 and Sec. 5 provide a comprehensive overview of these aspects for each paper studied in our survey. In general, diverse non-unified simulator choices result in task- and application-dependent evaluation protocols. As a result, it is difficult to compare the performance of different AIPP algorithms in a generalizable and fair fashion under consistent conditions.
@@
    Geometric approaches, which pre-compute paths to maximize coverage of the initially unknown environment for uniform data collection. Examples include lawnmower-like grid patterns (Blum et al., 2019; Westheider et al., 2023; Galceran and Carreras, 2013) or spiral- or circle-like patterns Zeng et al. (2022); Popović et al. (2020); Zhang et al. (2023).
@@
    Random walk-like methods, which sample paths at random often combined with heuristics to efficiently manage the mission budget (Zurbrügg et al., 2022; Dai et al., 2022; Kumar et al., 2022; Sukkar et al., 2019; Liu et al., 2023).
In addition to these simple non-adaptive benchmarks, many works consider variants of well-established planning algorithms in the AIPP context. Classical planners used for evaluation include branch-and-bound techniques (Binney and Sukhatme, 2012) and sampling-based methods, e.g., Monte Carlo Tree Search (MCTSChoudhury et al. (2020), rapidly exploring information gathering (RIG) trees Hollinger and Sukhatme (2014), or rapidly exploring random trees (Karaman and Frazzoli, 2011). Other benchmarks include geometric strategies, e.g., greedy frontier-based exploration (Yamauchi, 1997), or optimization-based routines, e.g., Bayesian optimization (Gelbart et al., 2014) or the covariance matrix adaptation evolution strategy (CMA-ES(Hansen and Ostermeier, 2001). To enable adaptive replanning during a mission as new sensor measurements arrive, these algorithms are commonly implemented in a fixed- or receding-horizon fashion.
@@ 4 Learning Approaches
This section surveys the AIPP literature from a learning-based perspective. We provide a taxonomy classifying relevant works based on the underlying learning algorithm they use, discuss pertinent aspects in each category, and integrate them in our unified mathematical framework for AIPP.
@@ 4.1 Supervised Learning
Traditionally, the components of the general AIPP framework in Fig. 1 are realized using hand-crafted models or heuristics, e.g., forward sensor models (Popović et al., 2020; Lim et al., 2016; Karaman and Frazzoli, 2011; Rückin et al., 2022a) or hand-crafted information criteria (Hollinger and Sukhatme, 2014; Binney and Sukhatme, 2012; Best et al., 2019; Isler et al., 2016). This makes it difficult to incorporate prior knowledge and the expected distributions of geometry and information in complex scenarios, as well as to achieve adaptability in new environments. To address these drawbacks, various methods have been proposed to learn different elements of mapping and planning through supervised training from labeled data.
@@
Supervised learning has been used for modeling environmental variables through Gaussian processes (Rayas Fernández et al., 2022; Cao et al., 2023a; Rückin et al., 2022a; Yang et al., 2023a; Cao et al., 2023c; Marchant et al., 2014; Hitz et al., 2017; Popović et al., 2020; Hollinger and Sukhatme, 2014; Vivaldini et al., 2019; Viseras and Garcia, 2019; Yanes Luis et al., 2023; Choi and Cielniak, 2021; Wei and Zheng, 2020; Song et al., 2015; Denniston et al., 2023b). As described in Sec. 3.1, Gaussian processes are often used in Eq. 2 due to their probabilistic and non-parametric nature, making them suitable for uncertainty-based information gathering with spatially complex underlying feature mappings $F$. In this context, supervised learning is used to compute the ground truth function $g=F$ from a training dataset $(x_{i},z_{i})\in\xi\times\mathcal{F}$, with $X=\xi$ and $Y=\mathcal{F}$, obtained by taking stochastic measurements $z_{i}$ in the environment. Note that we use $x_{i}$ to denote general training inputs and $\mathbf{x}_{i}$ to denote the robot pose. In Gaussian process mapping, each $x_{i}$ corresponds to the pose $\mathbf{x}_{i}\in\xi$ introduced in Sec. 2. Example applications include mapping distributions of temperature (Popović et al., 2020; Westheider et al., 2023; Rückin et al., 2022a), light (Denniston et al., 2023a; Cao et al., 2023a), water properties (Hitz et al., 2017; Yanes Luis et al., 2023), and vegetation (Popović et al., 2020; Vivaldini et al., 2019). The learning process in implicit neural representations can be interpreted similarly. These methods learn a function that maps 3D coordinates $X=\xi$ to radiance values $Y=\mathcal{F}$, which represent color and light intensity, from a dataset of posed RGB (Jin et al., 2023; Pan et al., 2022; Zhan et al., 2023; Sünderhauf et al., 2023; Pan et al., 2023b) or RGB-D (Ran et al., 2022; Yan et al., 2023; Lee et al., 2022; Zhang et al., 2023) images.
@@
For robotic exploration, several works focus on learning map completion from a partially observed map state, represented by its geometry (Shrestha et al., 2019; Zwecher et al., 2022; Georgakis et al., 2022a; Tao et al., 2023; Dhami et al., 2023; Li et al., 2023), topology (Saroya et al., 2020), or semantics (Schmid et al., 2023). At mission time step $t$, these methods predict the evolution of a probabilistic map representation $k$ steps into the future $\tilde{F}_{t+k}$. The dataset pairs $(x_{i},y_{i})$ used for learning map completion are application-dependent; for example, learning inputs can range from local observations (Schmid et al., 2023) to the current global map state (Li et al., 2020). The key advantage of such approaches is their ability to extrapolate incomplete maps by minimizing a specific loss during training. This extrapolation enables better reasoning with limited information and occlusions compared to classical heuristic strategies that plan solely based on the map state at a given time step.
@@
Several works have also investigate the benefits of using supervised learning in the planning algorithm. One straightforward approach is to learn the information criterion $I(\cdot)$ in Eq. 1 from a dataset of actions and their associated information gains observed in previous data gathering missions $(x_{i},y_{i})=(a_{i},I(a_{i}))$, where $X=\mathcal{A}$ and $Y=\mathbb{R}$ (Hepp et al., 2018; Bai et al., 2017; Ly and Tsai, 2019). To solve the AIPP problem in Eq. 1 directly, the learning process can depend on pairs that connect map states to corresponding actions. Each action $a_{i}$ is created using a planning algorithm with full access to the ground truth data represented by $g$. Schmid et al. (2022) and Zacchini et al. (2023) propose learning a probabilistic model of informative views in the context of sampling-based planning. Alternatively, the next best viewpoint can be learned by reasoning about the shape of an object to be reconstructed (Vasquez-Gomez et al., 2021; Mendoza et al., 2020) or localization performance in the context of active SLAM (Hanlon et al., 2023). When the dataset pairs $(x_{i},y_{i})$ are generated through expert demonstrations, supervised learning can be thought of as a simple form of imitation learning (IL) as described in Sec. 4.3. By learning the information criterion $I(\cdot)$ from data, these approaches all bypass formulating it explicitly for online replanning.
@@
A key requirement for all methods in Table 3 is the availability of reliable training data for model supervision. Some supervised learning methods in AIPP rely on open-source datasets, such as the indoor Matterport3D dataset (Georgakis et al., 2022a; Tao et al., 2023; Li et al., 2023) or outdoor 3D Street View dataset (Hepp et al., 2018). Works in environmental monitoring may use data from previously executed missions (Rayas Fernández et al., 2022; Hitz et al., 2017; Hollinger and Sukhatme, 2014; Caley et al., 2019), while others create custom synthetic environments approximating real-world scenarios (Schmid et al., 2022; Bai et al., 2017; Saroya et al., 2020; Cao et al., 2023a; Rückin et al., 2022a; Yang et al., 2023a; Hitz et al., 2017; Popović et al., 2020, 2020; Viseras and Garcia, 2019; Zwecher et al., 2022; Song et al., 2015; Marchant et al., 2014). However, the lack of realistic labeled training data for AIPP problems remains an open issue restricting the generalizability of existing methods to new environments and domains. Furthermore, since supervised learning models are trained only on static data, their applicability is limited in dynamic environments with changing obstacles or terrains or the target information distribution varies over time. Finally, as most deep learning models do not consider uncertainty in the learning predictions, they lack natural mechanisms to handle sensor noise or other sources of uncertainty. In Sec. 4.4, we discuss how active learning (AL) techniques can mitigate some of these issues.
@@
Architecture/method
Conditional variational autoencoder
(Schmid et al., 2022)
Convolutional neural network
(Schmid et al., 2022; Bai et al., 2017; Ly and Tsai, 2019; Caley and Hollinger, 2020; Caley et al., 2019; Zwecher et al., 2022; Hepp et al., 2018; Vasquez-Gomez et al., 2021; Mendoza et al., 2020; Li et al., 2023; Saroya et al., 2020; Georgakis et al., 2022a; Ramakrishnan et al., 2020; Li et al., 2023; Pan et al., 2023b, a)
Gaussian process
(Rayas Fernández et al., 2022; Cao et al., 2023a; Rückin et al., 2022a; Yang et al., 2023a; Cao et al., 2023c; Marchant et al., 2014; Hitz et al., 2017; Popović et al., 2020; Hollinger and Sukhatme, 2014; Vivaldini et al., 2019; Viseras and Garcia, 2019; Yanes Luis et al., 2023; Choi and Cielniak, 2021; Wei and Zheng, 2020; Song et al., 2015; Denniston et al., 2023b)
Implicit neural representation
(Jin et al., 2023; Pan et al., 2022; Ran et al., 2022; Zhan et al., 2023; Sünderhauf et al., 2023; Yan et al., 2023; Lee et al., 2022; Zhang et al., 2023; Pan et al., 2023b)
Kernel density estimation
(Zacchini et al., 2023)
Occupancy prediction network
(Tao et al., 2023)
Scene completion network
(Schmid et al., 2023)
Transformer
(Dhami et al., 2023; Hanlon et al., 2023)
Variational autoencoder
(Shrestha et al., 2019)
Training objective
3D semantic scene completion
(Schmid et al., 2023)
2D occupancy map completion
(Shrestha et al., 2019; Zwecher et al., 2022; Georgakis et al., 2022a; Ramakrishnan et al., 2020)
3D occupancy map completion
(Tao et al., 2023)
Environment topology
(Saroya et al., 2020)
Environmental variable
(Rayas Fernández et al., 2022; Cao et al., 2023a; Rückin et al., 2022a; Yang et al., 2023a; Cao et al., 2023c; Marchant et al., 2014; Hitz et al., 2017; Popović et al., 2020; Hollinger and Sukhatme, 2014; Vivaldini et al., 2019; Popović et al., 2020; Viseras and Garcia, 2019; Yanes Luis et al., 2023; Choi and Cielniak, 2021; Caley and Hollinger, 2020; Wei and Zheng, 2020; Song et al., 2015; Denniston et al., 2023b)
Grasp affordance
(Zhang et al., 2023)
Informative view distribution
(Schmid et al., 2022; Zacchini et al., 2023)
Localization score
(Hanlon et al., 2023)
Minimum view subset for coverage
(Pan et al., 2023a)
Next-best view
(Vasquez-Gomez et al., 2021; Mendoza et al., 2020)
Point cloud shape completion
(Dhami et al., 2023)
Radiance field
(Jin et al., 2023; Pan et al., 2022; Zhang et al., 2023; Ran et al., 2022; Zhan et al., 2023; Lee et al., 2022; Sünderhauf et al., 2023; Yan et al., 2023)
Required number of views
(Pan et al., 2023b)
Target point of interest
(Caley et al., 2019)
Unexplored navigable area beyond map frontiers
(Li et al., 2023)
Utility/reward
(Hepp et al., 2018; Schmid et al., 2022; Bai et al., 2017; Ly and Tsai, 2019)
Mission objective
Coverage
(Schmid et al., 2023, 2022; Saroya et al., 2020; Ly and Tsai, 2019; Vasquez-Gomez et al., 2021; Mendoza et al., 2020; Zwecher et al., 2022; Tao et al., 2023; Li et al., 2023)
Estimate quantile values
(Rayas Fernández et al., 2022)
Find maximum field value
(Song et al., 2015; Denniston et al., 2023b)
Maximize grasp quality
(Zhang et al., 2023)
Maximize localization accuracy
(Hanlon et al., 2023)
Maximize probability of improvement/expected improvement
(Denniston et al., 2023b)
Maximize observed surfaces
(Dhami et al., 2023; Pan et al., 2023a)
Minimize explored area to target point of interest
(Caley et al., 2019)
Minimize map error
(Viseras and Garcia, 2019; Choi and Cielniak, 2021; Caley and Hollinger, 2020; Ramakrishnan et al., 2020)
Minimize map uncertainty
(Hepp et al., 2018; Bai et al., 2017; Shrestha et al., 2019; Zacchini et al., 2023; Hollinger and Sukhatme, 2014; Yanes Luis et al., 2023; Georgakis et al., 2022a; Wei and Zheng, 2020; Denniston et al., 2023b)
Minimize map uncertainty and robot localization uncertainty
(Popović et al., 2020)
Minimize map uncertainty in high-interest areas
(Cao et al., 2023a; Rückin et al., 2022a; Yang et al., 2023a; Marchant et al., 2014; Hitz et al., 2017; Popović et al., 2020; Hollinger and Sukhatme, 2014)
Minimize map uncertainty in unclassified areas
(Vivaldini et al., 2019)
Minimize model uncertainty
(Jin et al., 2023; Pan et al., 2022; Ran et al., 2022; Zhan et al., 2023; Lee et al., 2022; Sünderhauf et al., 2023; Yan et al., 2023)
Minimize target localization uncertainty
(Cao et al., 2023c)
Object reconstruction based on PSNR
(Pan et al., 2023b)
Point-goal navigation
(Georgakis et al., 2022a; Ramakrishnan et al., 2020)
Training data
3D Street View dataset
(Hepp et al., 2018)
Custom CAD models
(Vasquez-Gomez et al., 2021; Mendoza et al., 2020; Ran et al., 2022; Yan et al., 2023; Pan et al., 2023a)
Custom synthetic dataset
(Schmid et al., 2022; Bai et al., 2017; Saroya et al., 2020; Cao et al., 2023a; Rückin et al., 2022a; Yang et al., 2023a; Hitz et al., 2017; Popović et al., 2020, 2020; Viseras and Garcia, 2019; Zwecher et al., 2022; Song et al., 2015; Marchant et al., 2014)
DTU dataset
(Jin et al., 2023)
Heuristic parameter setting
(Cao et al., 2023c; Yanes Luis et al., 2023)
Gibson dataset
(Ramakrishnan et al., 2020)
INRIA Aerial Image Labeling dataset
(Ly and Tsai, 2019)
KTH dataset
(Shrestha et al., 2019; Caley et al., 2019)
LLFF dataset
(Pan et al., 2022; Sünderhauf et al., 2023)
Matterport3D dataset
(Georgakis et al., 2022a; Tao et al., 2023; Li et al., 2023; Hanlon et al., 2023)
NeRF/Synthetic-NeRF dataset
(Pan et al., 2022; Ran et al., 2022; Lee et al., 2022; Sünderhauf et al., 2023; Zhan et al., 2023)
NYU dataset
(Schmid et al., 2023)
Real-world environmental data
(Rayas Fernández et al., 2022; Hitz et al., 2017; Hollinger and Sukhatme, 2014; Caley et al., 2019; Wei and Zheng, 2020; Denniston et al., 2023b; Vivaldini et al., 2019)
Realistic simulator
(Zacchini et al., 2023; Zhang et al., 2023)
Regional Ocean Modeling System (ROMS) dataset
(Caley and Hollinger, 2020)
ShapeNet dataset
(Dhami et al., 2023; Jin et al., 2023; Pan et al., 2023b)
Tanks&Temples dataset
(Zhan et al., 2023)

### v3

@@ 3.3 Benchmarks
3.3.1 Simulations
Simulators for AIPP evaluation make various interdependent design choices leading to highly specific simulation environments. First, AIPP methods are typically evaluated on a single fixed task, e.g., the RockSample task [62, 7], monitoring missions [3, 5, 27, 2, 69, 68], or reconstruction [43, 12, 48, 93, 57, 79, 80, 83]. Second, simulators differ in the robot configuration defined by the robot type, e.g., autonomous underwater vehicle (AUV[73], unmanned aerial vehicle (UAV[3], unmanned ground vehicle (UGV[66], or a robot arm [43], as well as the sensors, e.g., depth sensors [41, 46, 40, 56], thermal camera [5], point-based sensors [2, 65, 63, 71], or abstractions of sensor models interpreting raw sensor data [3, 92, 4, 94, 59]. Third, works vary in their assumptions about the structure of environments, e.g., 2D terrains [65, 3, 2, 5, 62, 7] or 3D volumes [41, 43, 1], and robot workspaces, e.g., obstacle-free 2D [2, 95], obstacle-free 3D [3, 27, 5, 59, 1], or obstacle-aware 3D [41, 43, 56] workspaces. Last, simulators vary according to the datasets they use, e.g., synthetic environments or randomly generated distributions [3, 2, 65, 5, 27, 39, 28, 96], more realistic simulators [43, 90, 45, 97], or real-world remote sensing datasets [3, 90, 27, 1, 65, 68, 49]. Sec. 4 and Sec. 5 provide a comprehensive overview of these aspects for each paper studied in our survey. In general, diverse non-unified simulator choices result in task- and application-dependent evaluation protocols. As a result, it is difficult to compare the performance of different AIPP algorithms in a generalizable and fair fashion under consistent conditions.
3.3.2 Common and Classical Baselines
@@
    Geometric approaches, which pre-compute paths to maximize coverage of the initially unknown environment for uniform data collection. Examples include lawnmower-like grid patterns [88, 27, 9] or spiral- or circle-like patterns [43, 3, 97].
@@
    Random walk-like methods, which sample paths at random often combined with heuristics to efficiently manage the mission budget [89, 55, 98, 30, 46].
In addition to these simple non-adaptive benchmarks, many works consider variants of well-established planning algorithms in the AIPP context. Classical planners used for evaluation include branch-and-bound techniques [99] and sampling-based methods, e.g., Monte Carlo Tree Search (MCTS[7], rapidly exploring information gathering (RIG) trees [68], or rapidly exploring random trees [100]. Other benchmarks include geometric strategies, e.g., greedy frontier-based exploration [101], or optimization-based routines, e.g., Bayesian optimization [102] or the covariance matrix adaptation evolution strategy (CMA-ES[103]. To enable adaptive replanning during a mission as new sensor measurements arrive, these algorithms are commonly implemented in a fixed- or receding-horizon fashion.
@@ 4 Learning Approaches
This section surveys the AIPP literature from a learning-based perspective. We provide a taxonomy classifying relevant works based on the underlying learning algorithm they use, discuss pertinent aspects in each category, and integrate them in our mathematical problem definition for AIPP.
@@ 4.1 Supervised Learning
Traditionally, the components of the general AIPP framework in Fig. 2 are realized using hand-crafted models or heuristics, e.g., forward sensor models [3, 104, 100, 5] or hand-crafted information criteria [68, 99, 75, 105]. This makes it difficult to incorporate prior knowledge and the expected distributions of geometry and information in complex scenarios, as well as to achieve adaptability in new environments. To address these drawbacks, various methods have been proposed to learn different elements of mapping and planning through supervised training from labeled data.
@@
Supervised learning has been used for modeling environmental variables through Gaussian processes [1, 2, 5, 63, 70, 71, 65, 3, 68, 67, 6, 64, 69, 66, 60, 59]. As described in Sec. 3.1, Gaussian processes are often used in Eq. 2 due to their probabilistic and non-parametric nature, making them suitable for uncertainty-based information gathering with spatially complex underlying feature mappings $F$. In this context, supervised learning is used to compute the ground truth function $g=F$ from a training dataset $(x_{i},z_{i})\in\xi\times\mathcal{F}$, with $X=\xi$ and $Y=\mathcal{F}$, obtained by taking stochastic measurements $z_{i}$ in the environment. Note that we use $x_{i}$ to denote general training inputs and $\mathbf{x}_{i}$ to denote the robot pose. In Gaussian process mapping, each $x_{i}$ corresponds to the pose $\mathbf{x}_{i}\in\xi$ introduced in Sec. 2. Example applications include mapping distributions of temperature [61, 27, 5], light [4, 2], water properties [65, 64], and vegetation [3, 67]. The learning process in implicit neural representations can be interpreted similarly. These methods learn a function that maps 3D coordinates $X=\xi$ to radiance values $Y=\mathcal{F}$, which represent color and light intensity, from a dataset of posed RGB [77, 78, 80, 81, 82] or RGB-D [79, 83, 84, 97] images.
@@
For robotic exploration, several works focus on learning map completion from a partially observed map state, represented by its geometry [44, 54, 36, 51, 57, 52], topology [96], or semantics [38]. At mission time step $t$, these methods predict the evolution of a probabilistic map representation $k$ steps into the future $\tilde{F}_{t+k}$. The dataset pairs $(x_{i},y_{i})$ used for learning map completion are application-dependent; for example, learning inputs can range from local observations [38] to the current global map state [106]. The key advantage of such approaches is their ability to extrapolate incomplete maps by minimizing a specific loss during training. This extrapolation enables better reasoning with limited information and occlusions compared to classical heuristic strategies that plan solely based on the map state at a given time step.
@@
Several works have also investigate the benefits of using supervised learning in the planning algorithm. One straightforward approach is to learn the information criterion $I(\cdot)$ in Eq. 1 from a dataset of actions and their associated information gains observed in previous data gathering missions $(x_{i},y_{i})=(a_{i},I(a_{i}))$, where $X=\mathcal{A}$ and $Y=\mathbb{R}$ [56, 28, 107]. To solve the AIPP problem in Eq. 1 directly, the learning process can depend on pairs that connect map states to corresponding actions. Each action $a_{i}$ is created using a planning algorithm with full access to the ground truth data represented by $g$. Schmid et al. [39] and Zacchini et al. [45] propose learning a probabilistic model of informative views in the context of sampling-based planning. Alternatively, the next best viewpoint can be learned by reasoning about the shape of an object to be reconstructed [47, 48] or localization performance in the context of active SLAM [108]. When the dataset pairs $(x_{i},y_{i})$ are generated through expert demonstrations, supervised learning can be thought of as a simple form of imitation learning (IL) as described in Sec. 4.3. By learning the information criterion $I(\cdot)$ from data, these approaches all bypass formulating it explicitly for online replanning.
@@
A key requirement for all methods in Table 3 is the availability of reliable training data for model supervision. Some supervised learning methods in AIPP rely on open-source datasets, such as the indoor Matterport3D dataset [36, 51, 52] or outdoor 3D Street View dataset [56]. Works in environmental monitoring may use data from previously executed missions [1, 65, 68, 49], while others create custom synthetic environments approximating real-world scenarios [39, 28, 96, 2, 5, 63, 65, 3, 61, 6, 54, 60, 71]. However, the lack of realistic labeled training data for AIPP problems remains an open issue restricting the generalizability of existing methods to new environments and domains. Furthermore, since supervised learning models are trained only on static data, their applicability is limited in dynamic environments with changing obstacles or terrains or the target information distribution varies over time. Finally, as most deep learning models do not consider uncertainty in the learning predictions, they lack natural mechanisms to handle sensor noise or other sources of uncertainty. In Sec. 4.4, we discuss how active learning (AL) techniques can mitigate some of these issues.
@@
Architecture/method Conditional variational autoencoder [39]
Convolutional neural network [39, 28, 107, 109, 49, 54, 56, 47, 48, 52, 96, 36, 37, 52, 82, 58]
Gaussian process [1, 2, 5, 63, 70, 71, 65, 3, 68, 67, 6, 64, 69, 66, 60, 59, 110, 111]
Implicit neural representation [77, 78, 79, 80, 81, 83, 84, 97, 82]
Kernel density estimation [45]
Occupancy prediction network [51]
Scene completion network [38]
Transformer [57, 108]
Variational autoencoder [44]
Training objective 3D semantic scene completion [38]
2D occupancy map completion [44, 54, 36, 37]
3D occupancy map completion [51]
Environment topology [96]
Environmental variable [1, 2, 5, 63, 70, 71, 65, 3, 68, 67, 61, 6, 64, 69, 109, 66, 60, 59, 110, 111]
Grasp affordance [97]
Informative view distribution [39, 45]
Localization score [108]
Minimum view subset for coverage [58]
Next-best view [47, 48]
Point cloud shape completion [57]
Radiance field [77, 78, 97, 79, 80, 84, 81, 83]
Required number of views [82]
Target point of interest [49]
Unexplored navigable area beyond map frontiers [52]
Utility/reward [56, 39, 28, 107]
Mission objective Coverage [38, 39, 96, 107, 47, 48, 54, 51, 52]
Estimate quantile values [1]
Find maximum field value [60, 59]
Maximize grasp quality [97]
Maximize localization accuracy [108]
Maximize probability of improvement/expected improvement [59, 110, 111]
Maximize observed surfaces [57, 58]
Minimize explored area to target point of interest [49]
Minimize map error [6, 69, 109, 37]
Minimize map uncertainty [56, 28, 44, 45, 68, 64, 36, 66, 59, 110, 111]
Minimize map uncertainty and robot localization uncertainty [61]
Minimize map uncertainty in high-interest areas [2, 5, 63, 71, 65, 3, 68]
Minimize map uncertainty in unclassified areas [67]
Minimize model uncertainty [77, 78, 79, 80, 84, 81, 83]
Minimize target localization uncertainty [70]
Object reconstruction based on PSNR [82]
Point-goal navigation [36, 37]
Training data 3D Street View dataset [56]
Custom CAD models [47, 48, 79, 83, 58]
Custom synthetic dataset [39, 28, 96, 2, 5, 63, 65, 3, 61, 6, 54, 60, 71]
DTU dataset [77]
Heuristic parameter setting [70, 64]
Gibson dataset [37]
INRIA Aerial Image Labeling dataset [107]
KTH dataset [44, 49]
LLFF dataset [78, 81]
Matterport3D dataset [36, 51, 52, 108]
NeRF/Synthetic-NeRF dataset [78, 79, 84, 81, 80]
NYU dataset [38]
Real-world environmental data [1, 65, 68, 49, 66, 59, 67]
Realistic simulator [45, 97]
Regional Ocean Modeling System (ROMS) dataset [109]
ShapeNet dataset [57, 77, 82]
Tanks&Temples dataset [80]